# Exact Sequence Interpolation with Transformers

### v2

@@ 5.6. Overall Construction and Complexity Estimate
Finally, we count the number of blocks and parameters in our transformer. The maximum width $d^{\prime}=4$ was required in the separation step, and at most
@@
blocks were needed in total. Moreover, the parameters in each block have been chosen as vectors in $\mathbb{R}^{d}$, rank-one matrices, multiples of the identity, or constants, resulting in $\mathcal{O}(d)$ parameters per block. The total number of non-zero parameters is therefore $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$.
6. Proof of Theorem˜1.6
We adapt the steps in Section˜5 to prove Theorem˜1.6. As discussed in Section˜1.3, the result is not simply obtained by continuous perturbation arguments, as these would only yield an approximate interpolation result. Indeed, a key limitation of the softmax self-attention formulation in ˜3.2 is that, for non-zero $A\in\mathbb{R}^{d\times d}$, even without residual ($\rho=0$) it always yields a strict combination of tokens, not allowing for exact collapse [^3] Setting $A=0$ and $\rho=0$, all tokens collapse exactly to their average, fundamentally limiting the output sequences length to $1$ (we point, however, to [12] for a workaround using $\rho=1$).. Therefore, we adapt the collapse step to include an extra transformer with $\mathcal{O}(N)$ blocks that utilizes non-residual feed-forward layers with ReLU activation to iteratively collapse small neighborhoods of tokens.
@@
6.1. Softmax Separation
The technical construction in Lemma˜5.1 needs to be adapted to ensure that there exists $\tau_{0}>0$ such that $\operatorname{\mathrm{T}}^{\tau}_{\text{sep}}(X^{j})\cap\operatorname{\mathrm{T}}^{\tau}_{\text{sep}}(X^{j^{\prime}})=\emptyset$ for all $j\neq j^{\prime}\in[N]$ and all $\tau\leq\tau_{0}$.
@@ 6.1. Softmax Separation
As discussed in Section˜3.2, for fixed $X$ and $A$, the softmax function $\pi^{\tau}$ and the hardmax function $\pi^{0}$ satisfy [9]
@@
(6.1) $\|\operatorname{\mathrm{SA}}^{0}_{i}(X)-\operatorname{\mathrm{SA}}^{\tau}_{i}(X)\|=|\alpha|\|\mathcal{M}_{i}^{0}(X,A)-\mathcal{M}_{i}^{\tau}(X,A)\|<|\alpha|\delta$
for all sufficiently small $\tau$. Moreover, every hardmax SA layer in the proof of Lemma˜5.1 ensures that tokens remain pairwise distinct. In particular, for some $\varepsilon>0$ and all $i\neq i^{\prime}\in[n]$ we have
@@
Using ˜6.1, we can similarly obtain for softmax SA layers that satisfy
@@
6.2. Softmax Leader Selection
We proceed exactly as in the leader selection step in the proof of Theorem˜1.4, with the extra requirement that there is “enough separation” between the desired positions of the $m^{j}$ tokens in each sequence that will become leaders in the next step. Denote $\hat{X}\coloneqq\operatorname{\mathrm{T}}_{\text{sep}}^{\tau}(X)$. For a given $\delta>0$, we choose $w\in\mathbb{R}^{d}$, $\|w\|=1$, and $R>0$ so the desired positions for the leaders in the next step $\{q_{i}^{j}\}_{i\in[m^{j}],j\in[N]}$ satisfy
(6.2) $\left\|w\left(\langle w,q_{i}^{j}-q_{i^{\prime}}^{j^{\prime}}\rangle\right)\right\|\geq 2\delta$
@@ 6.2. Softmax Leader Selection
(6.3) $\left\|w\left(\langle w,q_{i}^{j}-q_{i^{\prime}}^{j^{\prime}}\rangle\right)\right\|=R\left|\langle w,\hat{q}_{i}^{j}-\hat{q}_{i^{\prime}}^{j^{\prime}}\rangle\right|,$
@@
and choose $R>{2\delta}/{\zeta}$ to satisfy ˜6.2.
Figure 8. Geometric construction of the softmax collapse step for the proof of Theorem˜1.6. In the left panel, the result of the first substep, with all tokens in a ball of radius $\delta$ of the leaders. In the right panel, the $\operatorname{\mathrm{FF}}$ layer used in the second substep.
6.3. Softmax Collapse
@@
In the first substep, for sufficiently small $\tau>0$ and thanks to ˜6.2, a softmax SA layer with $\rho=0$, $V=I_{d}$ and $A=I_{d}$ moves all tokens of each sequence to a ball of radius $\delta$ of the $m^{j}$ leaders of that sequence, arranged in the previous leader selection step. The resulting configuration is shown in the left panel of Figure˜8 for a low-dimensional example.
@@
We relabel the tokens $\hat{x}_{1}^{1},\dots,\hat{x}_{1}^{N}$ if necessary, so that they are ordered from farthest ($\hat{x}_{1}^{1}$) to closest ($\hat{x}_{1}^{N}$) to the origin. Now, fix a non-residual ($\eta=0$) FF layer of width $d^{\prime}=3$ as follows (see right panel of Figure˜8):
@@
  • the remaining $m^{j}-1$ leaders per sequence $\hat{x}_{i}^{j}\in\mathbb{R}^{d}_{<0}$, $i\neq 1$, are projected to distinct (as ensured by ˜6.2) positive multiples of $w$, thus remaining in $\mathbb{R}^{d}_{<0}$.
@@
After $N$ iterations, we obtain a transformer $\operatorname{\mathrm{T}}^{\tau}_{\text{col}}$ of width $d^{\prime}=3$ and $N$ blocks, whose SA layers act as the identity ($\rho=1$, $V=0$, $A=0$). This transformer collapses each input sequence to exactly $m^{j}$ distinct tokens, after which the interpolation construction of Theorem˜1.4 applies verbatim.
@@
Our construction only requires that, within each self-attention layer, within each self-attention layer, $\tau$ be sufficiently small to achieve the desired behavior. This does not imply that subsequent layers produce identical outputs across models, as illustrated in Figure˜9:
@@
Figure 9. Illustration of Remark˜6.1 for $N=3$ sequences in $\mathbb{R}^{2}$. In panel $(a)$, the initial sequences’ configuration, with tokens from different sequences denoted with different colors and shapes. Panel $(b)$ shows the effect of one hardmax separation block, while panel $(c)$ shows the effect of a softmax block with $\tau=1$. Unlike the hardmax case, the softmax block can already separate all sequences, allowing the construction to proceed directly to the collapse step.
7. Conclusion
@@
7.1. Summary of Results
We presented the first exact interpolation result for transformers in sequence-to-sequence tasks with real vector-valued outputs. We constructed an explicit transformer architecture that performs exact sequence interpolation with a number of parameters independent of the input sequence length. Our construction mirrors key features of practical transformers, including the alternation of self-attention and feed-forward layers, and the low-rank structure of self-attention.
@@ 7.1. Summary of Results
7.2. Perspectives and Open Questions
@@
7.2.1. Complexity, Parameter Norms and Separation
While our transformer construction exhibits lower complexity, measured in terms of the number of blocks and parameters, than those proposed in previous works [12], we do not know whether it is optimal. In practice, transformers often achieve excellent approximate interpolation performance with a number of blocks significantly smaller than the dataset size. This raises the question of whether exact interpolation could also be attained with lower complexity. Addressing this question remains an interesting open problem.
Beyond block and parameter counts, an important open direction is to understand the magnitude of the parameters required in our construction. A priori bounds on the norm of the parameters would be relevant, for example, in the context of regularized training as pointed out in Section˜2. Intuitively, one expects the required norms to depend on how well-separated the tokens are in each step of our construction: well-separated configurations may be updated as desired using transformer blocks with moderate parameter values, whereas tightly clustered data may necessitate larger parameters. An illustrative example arises in the FF layers acting as “hat functions” of Lemma˜4.8, where the parameters need to blow up as data separation tends to zero. A systematic study of this effect, both theoretically and numerically, would provide valuable insight into exact interpolation constructions and is left as further work.
7.2.2. Extension to Probability Vector Outputs
@@
One can easily obtain an exactly interpolating transformer $\operatorname{\mathrm{T}}$ which outputs $Y$, provided $d=\mathcal{V}$. In practice, one typically works with an embedding dimension $d$ much smaller than $\mathcal{V}$. Extending our framework to this low-dimensional regime remains an interesting direction for future work.
7.2.3. Multi-Head Attention
Another relevant question is whether our arguments can be efficiently extended to transformers with multi-head attention. As heads can be seen as an analogue of width in ResNets, where a trade-off with depth has been shown [3], we wonder if multiple heads can reduce the number of parameters and blocks required to achieve perfect sequence interpolation. For this, one should replace ˜3.1 with
@@ 7.2.3. Multi-Head Attention
for a given number of heads $H\in\mathbb{N}$ and head-specific parameters $A^{h},V^{h}\in\mathbb{R}^{d\times d}$, $h\in[H]$. Our current analysis rigorously covers the case of a single head $H=1$ and justifies the good performance of these models given sufficient depth [24]. It would be interesting to investigate whether the additional degrees of freedom provided by multiple heads can reduce the required depth of the transformer, in particular the number of blocks used in the separation step described in Section˜5.2.
7.2.4. Masked Self-Attention
Masking is a common feature of self-attention layers for next-token interpolation tasks in natural language processing, where a token $x_{i}$ should only be influenced by those preceding it. To model this, one should replace ˜3.1 with
@@ 7.2.4. Masked Self-Attention
Notice that, for each $i\in[\operatorname{len}(X)]$, the sum runs only over $\ell\leq i$. This masked self-attention setting can be handled by our results in Theorem˜1.4 and Theorem˜1.6 through a problem reformulation as follows (we assume that $n^{j}=n$ and $m^{j}=m$ to simplify the exposition). For all $j\in[N]$,
@@
Now, we can apply Theorem˜1.4 or Theorem˜1.6 to the dataset $\{(X_{t}^{j},Y_{t}^{j})\}_{j\in[N],t\in[n]}$ consisting of $Nn$ sequences, and obtain a transformer $\operatorname{\mathrm{T}}$ such that $\operatorname{\mathrm{T}}(X^{j}_{t})=Y^{j}_{t}$ with $\mathcal{O}(Nn)$ blocks and $\mathcal{O}(dNn)$ parameters. Thus, it holds for a causal transformer $\operatorname{\mathrm{T}}_{\text{causal}}$ that
@@
as desired. This reformulation leads to a complexity that scales with input length $n$ rather than output length $m$, a scaling also observed in large-scale models [16]. Designing architectures that avoid this dependence while preserving exact interpolation remains an open challenge.
Acknowledgments
The authors thank Martín Hernández, Domènec Ruiz-Balet and Borjan Geshkovski for valuable conversations.
Funding. The work of AA was funded by the European Union’s Horizon Europe MSCA project ModConFlex (grant number 101073558). The work of EZ was funded by the Alexander von Humboldt-Professorship program, the ERC Advanced Grant CoDeFeL, the Grants PID2020-112617GB-C22 KiLearn and TED2021-131390B-I00-DasEl of MINECO and PID2023-146872OB-I00-DyCMaMod of MICIU (Spain), the European Union’s Horizon Europe MSCA project ModConFlex (grant number 101073558), the Transregio 154 Project “Mathematical Modelling, Simulation and Optimization Using the Example of Gas Networks” of the DFG, the AFOSR 24IOE027 project, and the Madrid Government–UAM Agreement for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation).

### v3

@@ 5.6 Overall construction and complexity estimate
Next, we count the number of blocks in our transformer. The maximum width $d^{\prime}=4$ was required in the separation step, and at most
@@
blocks were needed in total. More explicitly, ignoring layers that act as the identity, the construction is a sequence of applications of the key technical lemmas from Section 4. The separation step uses one FF layer to shift all tokens, followed by 2(N1)2(N-1) transformer blocks constructed with Lemma 5.1 to make the $N$ sequences pairwise disjoint. In the leader selection step, we shift all tokens to the positive orthant using one FF layer, then apply $\sum_{j=1}^{N}m^{j}$ FF layers of the type provided by Lemma 4.8 to position the leaders. The collapse step applies one self-attention layer of the type constructed in Lemma 4.4 to collapse tokens onto the leaders. Lastly, in the interpolation step we apply $\sum_{j=1}^{N}m^{j}$ FF layers provided by Lemma 4.8 to move the collapsed tokens to the desired outputs.
Finally, the parameters in each block have been chosen as vectors in $\mathbb{R}^{d}$, rank-one matrices, multiples of the identity, or constants, resulting in $\mathcal{O}(d)$ parameters per block. The total number of parameters is therefore $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$.
6 Proof of Theorem 1.8
We adapt the steps in Section 5 to prove Theorem 1.8. As discussed in Section 1.3, the result is not simply obtained by continuous perturbation arguments, as these would only yield an approximate interpolation result. Indeed, a key limitation of the softmax self-attention formulation in 3.2 is that, for non-zero $A\in\mathbb{R}^{d\times d}$, even without residual ($\rho=0$) it always yields a strict combination of tokens, not allowing for exact collapse [^2] Setting $A=0$ and $\rho=0$, all tokens collapse exactly to their average, fundamentally limiting the output sequences length to $1$ (we point, however, to [geshkovski2024measure] for a workaround using $\rho=1$).. Therefore, we adapt the collapse step to include an extra transformer with $\mathcal{O}(N)$ blocks that utilizes non-residual feed-forward layers with ReLU activation to iteratively collapse small neighborhoods of tokens.
@@
6.1 Softmax separation
The technical construction in Lemma 5.1 needs to be adapted to ensure that there exists $\tau_{0}>0$ such that $\operatorname{\mathrm{T}}^{\tau}_{\text{sep}}(X^{j})\cap\operatorname{\mathrm{T}}^{\tau}_{\text{sep}}(X^{j^{\prime}})=\emptyset$ for all $j\neq j^{\prime}\in[N]$ and all $\tau\leq\tau_{0}$.
@@ 6.1 Softmax separation
As discussed in Section 3.2, for fixed $X$ and $A$, the softmax function $\pi^{\tau}$ and the hardmax function $\pi^{0}$ satisfy [elfadelSoftmax1993]
@@
$\|\operatorname{\mathrm{SA}}^{0}_{i}(X)-\operatorname{\mathrm{SA}}^{\tau}_{i}(X)\|=|\alpha|\|\mathcal{M}_{i}^{0}(X,A)-\mathcal{M}_{i}^{\tau}(X,A)\|<|\alpha|\delta$ (6.1)
for all sufficiently small $\tau$. Moreover, every hardmax SA layer in the proof of Lemma 5.1 ensures that tokens remain pairwise distinct. In particular, for some $\varepsilon>0$ and all $i\neq i^{\prime}\in[n]$ we have
@@
Using 6.1, we can similarly obtain for softmax SA layers that satisfy
@@
6.2 Softmax leader selection
We proceed exactly as in the leader selection step in the proof of Theorem 1.6, with the extra requirement that there is “enough separation” between the desired positions of the $m^{j}$ tokens in each sequence that will become leaders in the next step. Denote $\hat{X}\coloneqq\operatorname{\mathrm{T}}_{\text{sep}}^{\tau}(X)$. For a given $\delta>0$, we choose $w\in\mathbb{R}^{d}$, $\|w\|=1$, and $R>0$ so the desired positions for the leaders in the next step $\{q_{i}^{j}\}_{i\in[m^{j}],j\in[N]}$ satisfy
$\left\|w\left(\langle w,q_{i}^{j}-q_{i^{\prime}}^{j^{\prime}}\rangle\right)\right\|\geq 2\delta$ (6.2)
@@ 6.2 Softmax leader selection
$\left\|w\left(\langle w,q_{i}^{j}-q_{i^{\prime}}^{j^{\prime}}\rangle\right)\right\|=R\left|\langle w,\hat{q}_{i}^{j}-\hat{q}_{i^{\prime}}^{j^{\prime}}\rangle\right|,$ (6.3)
@@
and choose $R>{2\delta}/{\zeta}$ to satisfy 6.2.
Figure 8: Geometric construction of the softmax collapse step for the proof of Theorem 1.8. In the left panel, the result of the first substep, with all tokens in a ball of radius $\delta$ of the leaders. In the right panel, the $\operatorname{\mathrm{FF}}$ layer used in the second substep.
6.3 Softmax collapse
@@
In the first substep, for sufficiently small $\tau>0$ and thanks to 6.2, a softmax SA layer with $\rho=0$, $V=I_{d}$ and $A=I_{d}$ moves all tokens of each sequence to a ball of radius $\delta$ of the $m^{j}$ leaders of that sequence, arranged in the previous leader selection step. The resulting configuration is shown in the left panel of Figure 8 for a low-dimensional example.
@@
We relabel the tokens $\hat{x}_{1}^{1},\dots,\hat{x}_{1}^{N}$ if necessary, so that they are ordered from farthest ($\hat{x}_{1}^{1}$) to closest ($\hat{x}_{1}^{N}$) to the origin. Now, fix a non-residual ($\eta=0$) FF layer of width $d^{\prime}=3$ as follows (see right panel of Figure 8):
@@
  • the remaining $m^{j}-1$ leaders per sequence $\hat{x}_{i}^{j}\in\mathbb{R}^{d}_{<0}$, $i\neq 1$, are projected to distinct (as ensured by 6.2) positive multiples of $w$, thus remaining in $\mathbb{R}^{d}_{<0}$.
@@
After $N$ iterations, we obtain a transformer $\operatorname{\mathrm{T}}^{\tau}_{\text{col}}$ of width $d^{\prime}=3$ and $N$ blocks, whose SA layers act as the identity ($\rho=1$, $V=0$, $A=0$). This transformer collapses each input sequence to exactly $m^{j}$ distinct tokens, after which the interpolation construction of Theorem 1.6 applies verbatim.
@@
Our construction only requires that, within each self-attention layer, $\tau$ be sufficiently small to achieve the desired behavior. This does not imply that subsequent layers produce identical outputs across models, as illustrated in Figure 9:
@@
Figure 9: Illustration of Remark 6.1 for $N=3$ sequences in $\mathbb{R}^{2}$. In panel $(a)$, the initial sequences’ configuration, with tokens from different sequences denoted with different colors and shapes. Panel $(b)$ shows the effect of one hardmax separation block, while panel $(c)$ shows the effect of a softmax block with $\tau=1$. Unlike the hardmax case, the softmax block can already separate all sequences, allowing the construction to proceed directly to the collapse step.
7 Conclusion
@@
7.1 Summary of results
We presented the first exact interpolation result for transformers in sequence-to-sequence tasks that applies to deep transformers with standard self-attention layers and to real vector-valued outputs. Specifically, we constructed an explicit deep transformer that exactly interpolates real-valued sequences with a number of parameters independent of the input sequence length. Our construction mirrors key features of practical transformers, including the alternation of self-attention and feed-forward layers, and the low-rank structure of self-attention.
@@ 7.1 Summary of results
7.2 Perspectives and open questions
@@
7.2.1 Complexity, parameter norms and separation
While our transformer construction exhibits lower complexity, measured in terms of the number of blocks and parameters, than those proposed in previous works [geshkovski2024measure], we do not know whether it is optimal. In practice, transformers often achieve excellent approximate interpolation performance with a number of blocks significantly smaller than the dataset size. This raises the question of whether exact interpolation could also be attained with lower complexity. Addressing this question remains an interesting open problem.
Beyond block and parameter counts, an important open direction is to understand the magnitude of the parameters required in our construction. A priori bounds on the norm of the parameters would be relevant, for example, in the context of regularized training as pointed out in Section 2. Intuitively, one expects the required norms to depend on how well-separated the tokens are in each step of our construction: well-separated configurations may be updated as desired using transformer blocks with moderate parameter values, whereas tightly clustered data may necessitate larger parameters. An illustrative example arises in the FF layers acting as “hat functions” of Lemma 4.8, where the parameters need to blow up as data separation tends to zero. A systematic study of this effect, both theoretically and numerically, would provide valuable insight into exact interpolation constructions and is left as further work.
7.2.2 Extension to probability vector outputs
@@
One can easily obtain an exactly interpolating transformer $\operatorname{\mathrm{T}}$ which outputs $Y$, provided $d=\mathcal{V}$. This probability-vector setting is closely related to the next-token prediction capacity studied in [madden2025next], where model complexity bounds are established for one-layer multi-head decoder-only transformers. The observation above shows that exact matching of finitely many probability vectors is also achievable within our construction when $d=\mathcal{V}$. In this sense, our result would be complementary to the results in [madden2025next], as it arises from a deep single-head sequence-to-sequence architecture. In practice, however, the embedding dimension $d$ is typically much smaller than $\mathcal{V}$. Extending our framework to this low-dimensional regime remains an interesting direction for future work.
7.2.3 Multi-head attention
Another relevant question is whether our arguments can be efficiently extended to transformers with multi-head attention. As heads can be seen as an analogue of width in ResNets, where a trade-off with depth has been shown [alvarez2024interplay], we wonder if multiple heads can reduce the number of parameters and blocks required to achieve perfect sequence interpolation. For this, one should replace 3.1 with
@@ 7.2.3 Multi-head attention
for a given number of heads $H\in\mathbb{N}$ and head-specific parameters $A^{h},V^{h}\in\mathbb{R}^{d\times d}$, $h\in[H]$. Our current analysis rigorously covers the case of a single head $H=1$ and justifies the good performance of these models given sufficient depth [michel2019sixteen]. It would be interesting to investigate whether the additional degrees of freedom provided by multiple heads can reduce the required depth of the transformer, in particular the number of blocks used in the separation step described in Section 5.2.
7.2.4 Masked self-attention
Masking is a common feature of self-attention layers for next-token interpolation tasks in natural language processing, where a token $x_{i}$ should only be influenced by those preceding it. To model this, one should replace 3.1 with
@@ 7.2.4 Masked self-attention
Notice that, for each $i\in[\operatorname{len}(X)]$, the sum runs only over $\ell\leq i$. This masked self-attention setting can be handled by our results in Theorem 1.6 and Theorem 1.8 through a problem reformulation as follows (we assume that $n^{j}=n$ and $m^{j}=m$ to simplify the exposition). For all $j\in[N]$,
@@
Now, we can apply Theorem 1.6 or Theorem 1.8 to the dataset $\{(X_{t}^{j},Y_{t}^{j})\}_{j\in[N],t\in[n]}$ consisting of $Nn$ sequences, and obtain a transformer $\operatorname{\mathrm{T}}$ such that $\operatorname{\mathrm{T}}(X^{j}_{t})=Y^{j}_{t}$ with $\mathcal{O}(Nn)$ blocks and $\mathcal{O}(dNn)$ parameters. Thus, it holds for a causal transformer $\operatorname{\mathrm{T}}_{\text{causal}}$ that
@@
as desired. This reformulation leads to a complexity that scales with input length $n$ rather than output length $m$, a scaling also observed in large-scale models [hoffmann2022training]. Designing architectures that avoid this dependence while preserving exact interpolation remains an open challenge.
[…]
Acknowledgements The authors thank Martín Hernández, Domènec Ruiz-Balet and Borjan Geshkovski for valuable conversations.
Declarations
  • Funding
    The work of AA was funded by the European Union’s Horizon Europe MSCA project ModConFlex (grant number 101073558). The work of GF is funded by the ANR-DFG project MONET (DFG project number 568735456). The work of EZ was funded by the Alexander von Humboldt-Professorship program, the ERC Advanced Grant CoDeFeL, the Grants PID2020-112617GB-C22 KiLearn and TED2021-131390B-I00-DasEl of MINECO and PID2023-146872OB-I00-DyCMaMod of MICIU (Spain), the European Union’s Horizon Europe MSCA project ModConFlex (grant number 101073558), the Transregio 154 Project “Mathematical Modelling, Simulation and Optimization Using the Example of Gas Networks” of the DFG, the AFOSR 24IOE027 project, the Madrid Government–UAM Agreement for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation), and the SURE-AI Centre grant 357482, Research Council of Norway.
  • Competing interests
    The authors declare no competing interests.
  • Ethics approval and consent to participate
    Not applicable
  • Consent for publication
    All authors consent to the publication of this manuscript.
  • Data availability
    Not applicable
  • Materials availability
    Not applicable
  • Code availability
    Code available in the GitHub repository:
  • Author contribution
    All authors contributed to the study conception and design. The initial idea was developed by A. Alcalde and E. Zuazua. The analysis was mainly performed by A. Alcalde and G. Fantuzzi. The manuscript was first written by A. Alcalde and subsequently reviewed and edited by all authors. All authors read and approved the final manuscript.